How to solve linear systems in 4 varibles.
for example:
W + X +Y + Z = 3
W - X + Y + Z = 1
W + X - Y + Z = 1
W + X + Y - Z = 3
___________________

Answer by thechamp1011(19)

W+X+Y+Z=3
W-X+Y+Z=1
W+X-Y+Z=1
W+X+Y-Z=3
___________________
4W+2X+2Y+2Z=8
2(2W+X+Y+Z)=8
2W+X+Y+Z=8/2=4